11. xºy(3) – 3x²y + xy = 0 (use a substitution v In x).

Find a solution
Answer #1

Solution e x3y _ 3x²y + xy=0 Let 1- Ina 2= ele y = dy dy zu ax e dy da = 3732 du do dus ay = dy dio y (w.o.tv zzy - dzm(m2-6m + 2) = 0 >m=0, 303 • Solution is y= creou (3+3) +cze +63 فاز ول۔ Sulestitute back - Inx Aus y = 4+ (223+53 3-3 +63 x